The length of a rectangle is 6 times the width. The area of the rectangle is 84cm^2.
Answer by mananth(16949)

let width be x
length = 6x
Area = L*W
6x^2=84
/6
x^2=14
x=3.74
width = 3.74
length = 6*3.74
length =22.45